# Polarization Diagnostics Applied to Coronal Mass Ejections and the Background Solar Wind

### v1

@@ 1 Introduction
A primary science target for PUNCH is characterizing Coronal Mass Ejections (CMEs; see, e.g., Malanushenko et al. (2025)) which are generally expected to incorporate coherently twisted magnetic fields, i.e., magnetic flux ropes (see, e.g., Provornikova et al. (2025)). PUNCH provides a full-circle field of view (FOV) from five solar radii outward, mosaicking the fields of view of a central Near Field Imager (NFI) (Colaninno et al., 2025) with three Wide Field Imagers (WFI) (Laurent, 2025). In this way, PUNCH fills existing gaps between coronagraphs and heliospheric imagers and obtains continuous wide-field observations of the evolution of CMEs and their substructure from the Sun throughout the inner solar system. This large FOV is also intended to yield new insight into the formation, evolution and 3D structure of stream interaction regions (SIRs) that arise between fast and slow wind streams (de Koning et al., 2025).
PUNCH obtains both brightness and polarized brightness measurements to image and locate evolving SIR and CME structures in 3D: a measurement of great importance to space weather predictions (Seaton et al., 2025). Until PUNCH, only coronagraphs (i.e., no heliospheric imagers) have had both brightness and polarized brightness observations. PUNCH is the first heliospheric imager to be designed with polarization capability and to have the sensitivity, spatial FOV, and polarization capability needed to locate and track bright structures as they propagate, with both its inner coronagraph and its three heliospheric imagers operating as a single “virtual instrument”.
@@
In Section 3, we present a variety of forward modeled CMEs, comparing the 3D position of the CME front to ground truth simulation knowledge of the position of mass along the line of sight. We also discuss how front-back ambiguity associated with polarization analysis can be resolved using the wide field of view and sensitivity of PUNCH images, and how establishing the 3D postion of CME substructures might be used to diagnose magnetic flux rope chirality.
@@ 2.1 Background
The brightness and polarization of the white light corona was first considered theoretically by Schuster (1879), and then in comparison to observations of solar eclipses (Young, 1911; Minnaert, 1930; Baumbach, 1938; Öhman, 1947; van de Hulst, 1950). These analyses demonstrated how radiation from the solar photosphere is Thomson scattered in the optically-thin corona, and how the linearly polarized component of that radiation (polarized brightness, or $pB$) becomes an excellent diagnostic of the line-of-sight integral of $n_{e}$ (see also Altschuler and Perry, 1972; Munro and Jackson, 1977; Inhester, 2015). The fall-off with distance of $n_{e}(r)$ has been studied using $pB$ by multiple authors, including Saito et al. (1970); Munro and Jackson (1977); Fisher and Guhathakurta (1995); Guhathakurta, Holzer, and MacQueen (1996); Gibson et al. (1999); Edwards et al. (2023). Others have combined coronagraph data with radio and in-situ measurements to find density parameterizations that cover distances from the low corona to interplanetary space (e.g., Bird and Edenhofer, 1990; Leblanc, Dulk, and Bougeret, 1998; Mann et al., 2023).
@@
We note that at least two definitions of $pB$ have entered the literature, as more fully detailed in DeForest, Seaton, and West (2022). In this paper, we consider $pB$ to be the difference in transmitted radiance through two ideal linear polarizers oriented tangentially and radially to the Sun, respectively, on an instrument image plane. In other words, $pB$ is a modified Stokes parameter analogous to Stokes $Q$ but with reference angle rotating according to the image plane geometry rather than fixed in the image plane.
Finally, we note that measurements of the white-light corona involve the combination of the electron-scattered K-corona and the dust-scattered F-corona (the latter being the cause of “zodiacal light”). The F-corona is essentially unpolarized in the inner corona (elongation angles $\varepsilon<5-10^{\circ}$), and so it does not contribute significantly to $pB$ there (Lamy et al., 2020). However, the F-corona does become polarized to approximately 7% at larger distances (Morgan and Cook, 2020). Further, total brightness ($tB$) includes unpolarized light and so has a contribution from the F-corona at all heights. Fortunately, the F-corona is very smooth compared to the K-corona, and robust methods have been established to model and remove it from both $pB$ and $tB$ (Morgan and Cook, 2020; Lamy et al., 2021; Lamy, Gilardy, and Llebaria, 2022).
@@ 2.2 Locating Position of Superparticles
Studies of the corona close to the Sun are able to use the approximation that the set of line-of-sight observing rays are all parallel to one another. In that limit, one can also define the “plane of the sky” (POS) as the plane perpendicular to those rays that also contains the center of the Sun. However, for large elongation angles (as observed by PUNCH’s Wide Field Imagers) this limit no longer applies and it makes more sense to refer to the Thomson Surface (TS; Figure 1), sometimes called the Thomson Sphere, as the locus of points that generally maximize the optically thin scattering signal (see, e.g., Vourlidas and Howard, 2006; Howard and DeForest, 2012; Howard et al., 2013; DeForest, Howard, and Tappin, 2013a).
Bearing this in mind, we now consider a compact mass, or superparticle, in 3D Cartesian coordinates, as shown in Figure 1. (We will explicitly define conditions where a structure’s central position can be found, such that it may be considered “localized”, in Section 2.5.) The line-of-sight elongation angle $\varepsilon$ gives the radial distance $r_{pos}$ projected on the YZY-Z plane, and this along with the object’s projected position angle $\alpha$ allows us to calculate a projected (or, for small $\varepsilon$, actual) Y and Z position, as shown in the bottom left insert of Figure 1. To find the 3D position we then need to determine the object’s position $x$ along the $X$ axis. As we show below, the degree of polarization, $p\equiv{pB}/{tB}$, can be directly related to the scattering angle $\chi$ of light off of the object. $\chi$ is the complement of $\tau$, the angular position of the object along the line of sight from the TS, and $x$ can be determined from either $\chi$ or $\tau$, in combination with $\varepsilon$ (see Section 2.2.2). The full 3D Cartesian position, $x,y,z$ is thus established.
This property of polarization has been exploited in previous works, (e.g., Hildner et al., 1975; Poland and Munro, 1976; Crifo, Picat, and Cailloux, 1983; Moran and Davila, 2004; Dere, Wang, and Howard, 2005; Moran, Davila, and Thompson, 2010; Mierla et al., 2009, 2010, 2011; Dai et al., 2014; Chifu, 2015; Pagano, Bemporad, and Mackay, 2015; Bemporad and Pagano, 2015; Lu et al., 2017; Floyd and Lamy, 2019; Dai, Wang, and Inhester, 2020; Mierla et al., 2022) One example that we will return to in Section 3.3 is DeForest, de Koning, and Elliott (2017a), who used polarization data in STEREO-A COR2 coronagraph data to map the position of the front and core of a CME in three dimensions. When they reprojected it onto the plane of sky observed by the SOHO LASCO coronagraph, they found consistent cotemporal structures between the two longitudinally displaced satellites.
@@ 2.2.1 Center of Mass
where $s$ is the distance along the line of sight from the Thomson Sphere (TS) to the scattering point. From Figure 1 we see that,
@@
where $d$ is the radius of closest approach of the line of sight, i.e., the intersection with the TS, and $\tau$ is the angle along the line of sight from this to each scattering point. Note that $d$ can be directly related to the line-of-sight elongation angle $\varepsilon$ and the observer’s distance from the Sun. Taking observations from Earth, we have:
@@ 2.2.2 Polarization Ratio
we can express this in terms of line-of-sight angle, $\tau$,
@@
which matches the far-field formula derived by DeForest, Howard, and Tappin (2013b).
@@
In summary: if sufficiently compact (more on this in Section 2.5), we can directly relate PUNCH observations of $pB$, tBtB to the 3D COM position of a superparticle — subject to a sign ambiguity as to whether it is in front or behind the TS (see Section 3.2). Once this ambiguity is resolved, the 3D position of the superparticle is obtained.
@@ 2.3 Diagnosing Radial Density Falloff
The density falloff in the background corona and solar wind is generally radial, i.e., $n_{e}\propto r^{-c}$ (Saito et al., 1970), with density exponent $c$ decreasing from values of order 10 to 20 in the very low corona (Gibson et al., 1999; Guhathakurta et al., 1999) where wind flow is subsonic and scale height effects dominate, to the asymptotic value of 2 expected for a constant-velocity solar wind in interplanetary space (Parker, 1957). We note that the fast solar wind is expected to reach its final speed closer to the Sun than the slow solar wind, which continues to accelerate to large distances so that $c>2$ (Bunting et al., 2024). For the range of distances relevant to PUNCH that we primarily consider in this paper, the expected range of variation in $c$ is between 22 and $4$, e.g., as shown in the observationally-constrained model of Leblanc, Dulk, and Bougeret (1998), $c\approx 3.35$ at $r=5\,R_{\odot}$, and it decreases to 2.30 at $r=10\,R_{\odot}$, then to 2.002 at $r=100\,R_{\odot}$. Using data from Parker Solar Probe, Maruca et al. (2023) found evidence for solar wind acceleration between 30 and 150 $R_{\odot}$, with an associated density exponent $c\approx 2.26$ in that region. One-dimensional turbulence-driven models provide similar ranges, with the full set of ZEPHYR models from Cranmer, van Ballegooijen, and Edgar (2007) and Cranmer, van Ballegooijen, and Woolsey (2013) having median values of $c$ of 2.97, 2.47, and 2.07 at $r=5,10$, and 100 $R_{\odot}$, respectively.
@@ 2.3.1 Center of Mass
Consider a spherically symmetric density falloff,
@@
we find for each line of sight of a given elongation angle, $\varepsilon$,
@@ 2.3.2 Polarization Ratio
As already noted, we can simplify by assuming the scattering occurs far enough from the Sun for it to be considered a point source (true for the PUNCH FOV $5R_{\odot}$). In that limit of large $r$, it follows from Equations 8-9 and 16 that
@@
Figure 2: Diagnostics obtained for a density distribution $n_{e}(r)=n_{o}\cdot r^{-2}$. Top: Ratio of white-light polarized brightness to total brightness ($p$), middle: polarization ratio (PRPR), and bottom: $\tau_{pol}$, all plotted vs elongation angle $\varepsilon$. The blue dash-dotted line shows the asymptotically regained symmetric solutions, $p_{symm}=3/5,PR_{symm}=1/4$, and $\tau_{pol_{symm}}=\operatorname{asin}(\sqrt{(}1/4))$. Differences are seen between this simplified point-source solution and the solid-black-line complete-form solution, which takes into account the extended disk of the Sun and truncates lines of sight at the observer, as discussed in the text.
@@
Figure 3: Forward modeled $\tau_{pol}$ for $n_{e}=r^{-c}$ spherically symmetric density distributions, for choices of (top) $c=1.9,2.1,2.2$ and (bottom) $c=2.5,3,4$, for elongation angles $\varepsilon<45^{\circ}$.
To establish how generally this slope diagnostic might be used, we need to consider the effects of asymmetric integration limits for lines of sight $\varepsilon>0$. In these cases, we see from Equation 2.3.2 that an additional correction term dependent on line-of-sight elongation angle $\varepsilon$ needs to be added:
@@
It is clear that as $\varepsilon\rightarrow 0$, $F_{corr}(\varepsilon)\rightarrow 0$, and similarly as $\varepsilon\rightarrow\frac{\pi}{2}$, $F_{corr}(\varepsilon)\rightarrow 0$. For both of these limits, the symmetric solutions Equations 38-40 are regained.
@@
Something close to a $r^{-2}$ falloff is what we would expect to see for a mass-conserving constant-speed solar wind, and indeed, Figure 4 shows $\tau_{pol_{r^{-2}}}=\operatorname{asin}(\sqrt{(}1/4))=0.524$ is the central value from a global simulation of a forward-modeled complex solar wind (also seen in Figures 2 and 3).
In summary: For observations along lines of sights with elongation angles in the range $\sim 1-15^{\circ}$, PUNCH white light polarization observations will provide a local diagnostic of density radial falloff for the background solar wind, independent from and complementary to non-local techniques. The effects of asymmetry are predictable and small enough to distinguish between falloff slopes in the lower part of the PUNCH/NFI field of view ($2-10r_{\odot}$ [^1] Note throughout this paper we use the $r_{\odot}$ to signify the fixed solar radius, $r_{\odot}=695700~km$. This is not to be confused with the unit $R_{\odot}$ sometimes used in DeForest et al. (2025) and other PUNCH papers, which is a convenience unit defined as the approximate apparent solar radius in elongation angle $\varepsilon$ when viewed from $1au$, i.e., $R_{\odot}=0.25^{\circ}$.).

### v2

@@ 1 Introduction
As discussed in other papers in this topical issue (Malanushenko et al., 2025; Provornikova et al., 2025), a primary science target for PUNCH is characterizing Coronal Mass Ejections (CMEs). PUNCH provides a full-circle field of view (FOV) from $1.5-45^{\circ}$ elongation or $6-180~R_{\odot}$ [^1] Because elongation angle $\varepsilon$ does not generally map linearly to solar radius, we use $R_{\odot}$ as a convenience unit representing the approximate apparent solar radius in elongation angle $\varepsilon$ when viewed from $1au$, i.e., $R_{\odot}=0.25^{\circ}$. It is not to be confused with $r_{\odot}$, used elsewhere in this paper, which signifies the physical and fixed solar radius, $r_{\odot}=695700~km$., mosaicking the fields of view of a central Near Field Imager (NFI) (Colaninno et al., 2025) with three Wide Field Imagers (WFI) (Laurent, 2025). In this way, PUNCH fills gaps between existing coronagraphs and heliospheric imagers (Brueckner et al., 1995; Howard et al., 2008; Tomczyk et al., 2016; Vourlidas et al., 2016; Howard et al., 2020; Zhukov et al., 2025b) and obtains continuous wide-field observations of the evolution of CMEs and their substructure from the Sun throughout the inner solar system. This large FOV also yields new insight into the formation, evolution and 3D structure of the stream interaction regions (SIRs) that arise between fast and slow wind streams (Neugebauer and Snyder, 1966; Rouillard et al., 2008) as discussed in this topical issue by de Koning et al. (2025).
PUNCH obtains both brightness and polarized brightness measurements to image and locate evolving SIR and CME structures in 3D: a measurement of great importance to space weather predictions (Seaton et al., 2025). Other coronagraphs including SOHO/LASCO (Moran and Davila, 2004) and STEREO/COR1 (Moran et al., 2010) have had both brightness and polarized brightness observations, but no heliospheric imagers until now. PUNCH is the first heliospheric imager to be designed with the sensitivity, spatial FOV, and polarization capability needed to locate and track bright structures as they propagate, with both its inner coronagraph and its three heliospheric imagers operating as a single “virtual instrument” (DeForest et al., 2025).
@@
In Section 3, we present a variety of forward modeled CMEs, comparing the 3D position of the CME front to ground truth simulation knowledge of the position of mass along the line of sight. We also discuss how front-back ambiguity associated with polarization analysis can be resolved using the wide field of view and sensitivity of PUNCH images, and how establishing the 3D position of CME substructures might be used to diagnose magnetic flux rope chirality.
@@ 2.1 Background
The brightness and polarization of the white light corona was first considered theoretically by Schuster (1879), and then in comparison to observations of solar eclipses (Young, 1911; Minnaert, 1930; Baumbach, 1938; Öhman, 1947; van de Hulst, 1950). These analyses demonstrated how radiation from the solar photosphere is Thomson-scattered in the optically-thin corona, and how the linearly polarized component of that radiation (polarized brightness, or $pB$) becomes an excellent diagnostic of the line-of-sight (LOS) integral of $n_{e}$ (see also Altschuler and Perry, 1972; Munro and Jackson, 1977; Inhester, 2015). The falloff with distance of $n_{e}(r)$ has been studied using $pB$ by multiple authors, including Saito et al. (1970); Munro and Jackson (1977); Fisher and Guhathakurta (1995); Guhathakurta et al. (1996); Gibson et al. (1999a); Edwards et al. (2023). Others have combined coronagraph data with radio and in-situ measurements to find density parameterizations that cover distances from the low corona to interplanetary space (e.g., Bird and Edenhofer, 1990; Leblanc et al., 1998; Mann et al., 2023).
@@
We note that at least two definitions of $pB$ have entered the literature, as more fully detailed in DeForest et al. (2022). In this paper, we consider $pB$ to be the difference in transmitted radiance through two ideal linear polarizers oriented tangentially and radially to the Sun, respectively, on an instrument image plane. In other words, $pB$ is a modified Stokes parameter analogous to Stokes $Q$ but with reference angle rotating according to the image plane geometry rather than fixed in the image plane (see DeForest et al. (2025) Appendix 2 for further discussion).
Finally, we note that measurements of the white-light corona involve the combination of the electron-scattered K-corona and the dust-scattered F-corona (the latter being the cause of “zodiacal light”). The F-corona is essentially unpolarized in the inner corona (elongation angles $\varepsilon<5-10^{\circ}$), and so it does not contribute significantly to $pB$ there (Lamy et al., 2020). However, the F-corona does become polarized to approximately 7% at larger distances (Morgan and Cook, 2020). Further, total brightness ($tB$) includes unpolarized light and so has a contribution from the F-corona at all heights. Fortunately, the F-corona is very smooth compared to the K-corona, and robust methods have been established to model and remove it from both $pB$ and $tB$ (Morgan and Cook, 2020; Lamy et al., 2021, 2022).
@@ 2.2 Locating Position of Superparticles
Studies of the corona close to the Sun are able to use the approximation that the set of LOS observing rays are all parallel to one another. In that limit, one can also define the “plane of the sky” (POS) as the plane perpendicular to those rays that also contains the center of the Sun. However, for large elongation angles (as observed by PUNCH’s Wide Field Imagers) this limit no longer applies and it makes more sense to refer to the Thomson Surface (TS; Figure 1), sometimes called the Thomson Sphere, as the locus of points that generally maximize the optically thin scattering signal (see, e.g., Vourlidas and Howard, 2006; Howard and DeForest, 2012; Howard et al., 2013; DeForest et al., 2013a).
Bearing this in mind, we now consider a compact mass, or superparticle, in 3D Cartesian coordinates, as shown in Figure 1. (We will explicitly define conditions where a structure’s central position can be found, such that it may be considered “localized”, in Section 2.5.) The LOS elongation angle $\varepsilon$ gives the radial distance $r_{pos}$ projected on the YZY-Z plane, and this along with the object’s projected position angle $\alpha$ allows us to calculate a projected (or, for small $\varepsilon$, actual) Y and Z position, as shown in the bottom left insert of Figure 1. To find the 3D position we then need to determine the object’s position $x$ along the $X$ axis. As we show below, the degree of polarization, $p\equiv{pB}/{tB}$, can be directly related to the scattering angle $\chi$ of light off of the object. $\chi$ is the complement of $\tau$, the angular position of the object along the line of sight from the TS, and $x$ can be determined from either $\chi$ or $\tau$, in combination with $\varepsilon$ (see Section 2.2.2). The full 3D Cartesian position, $x,y,z$ is thus established.
This property of polarization has been exploited in previous works, (e.g., Hildner et al., 1975; Poland and Munro, 1976; Crifo et al., 1983; Moran and Davila, 2004; Dere et al., 2005; Moran et al., 2010; Mierla et al., 2009, 2010, 2011; Dai et al., 2014; Chifu, 2015; Pagano et al., 2015; Bemporad and Pagano, 2015; Lu et al., 2017; Floyd and Lamy, 2019; Dai et al., 2020; Mierla et al., 2022). One example that we will return to in Section 3.3 is DeForest et al. (2017b), who used polarization data in STEREO-A COR2 coronagraph data to map the position of the front and core of a CME in three dimensions. When they reprojected it onto the plane of sky observed by the SOHO LASCO coronagraph, they found consistent cotemporal structures between the two longitudinally displaced satellites.
@@ 2.2.1 Center of Mass
where $s$ is the distance along the line of sight from the TS to the scattering point. From Figure 1 we see that,
@@
where $d$ is the distance of closest approach of the line of sight, i.e., the intersection with the TS, and $\tau$ is the angle along the line of sight from this to each scattering point. Note that $d$ can be directly related to the LOS elongation angle $\varepsilon$ and the observer’s distance from the Sun. Taking observations from Earth, we have:
@@ 2.2.2 Polarization Ratio
we can express this in terms of the angle $\tau$,
@@
which matches the far-field formula derived by DeForest et al. (2013b).
@@
In summary: if sufficiently compact (more on this in Section 2.5), we can directly relate PUNCH observations of $pB$, tBtB to the 3D COM position of a superparticle — subject to a sign ambiguity as to whether it is in front or behind the TS. Once this ambiguity is resolved by polarization data (see Section 3.2), the 3D position of the superparticle is obtained.
@@ 2.3 Diagnosing Radial Density Falloff
The density falloff in the background corona and solar wind is generally radial, i.e., $n_{e}\propto r^{-c}$ (Saito et al., 1970), with density exponent $c$ decreasing from values of order 10 in the first few tenths of a solar radius (Gibson et al., 1999b; Guhathakurta et al., 1999) where wind flow is subsonic and scale height effects dominate, to the asymptotic value of 2 expected for a constant-velocity solar wind in interplanetary space (Parker, 1957). We note that the fast solar wind is expected to reach its final speed closer to the Sun than the slow solar wind, which continues to accelerate to large distances so that $c>2$ (Bunting et al., 2024). For the range of distances relevant to PUNCH that we primarily consider in this paper, the expected range of variation in $c$ is between 22 and $4$, e.g., as shown in the observationally-constrained model of Leblanc et al. (1998), $c\approx 3.35$ at $r=5\,R_{\odot}$, and it decreases to 2.30 at $r=10\,R_{\odot}$, then to 2.002 at $r=100\,R_{\odot}$. Using data from Parker Solar Probe, Maruca et al. (2023) found evidence for solar wind acceleration between 30 and 150 $R_{\odot}$, with an associated density exponent $c\approx 2.26$ in that region. One-dimensional turbulence-driven models provide similar ranges, with the full set of models from Cranmer et al. (2007) and Cranmer et al. (2013) having median values of $c$ of 2.97, 2.47, and 2.07 at $r=5,10$, and 100 $R_{\odot}$, respectively.
@@ 2.3.1 Center of Mass
Consider a spherically symmetric density falloff with distance,
@@
we find for each line of sight of a given elongation angle, $\varepsilon$, and limits of integration along the line of sight $\tau_{-}$ to $\tau_{+}$,
@@ 2.3.2 Polarization Ratio
As already noted, we can simplify by assuming the scattering occurs far enough from the Sun for it to be considered a point source (true for the PUNCH FOV $5R_{\odot}$). In that limit of large $r$ where the limb darkening term drops out, it follows from Equations 8-9 and 16 that
@@
Figure 2: Diagnostics obtained for a density distribution $n_{e}(r)=n_{o}\cdot r^{-2}$. Top: Ratio of white-light polarized brightness to total brightness ($p$), middle: polarization ratio (PRPR), and bottom: $\tau_{pol}$, all plotted vs. elongation angle $\varepsilon$. The blue dash-dotted line shows the asymptotically regained symmetric solutions, $p_{symm}=3/5,PR_{symm}=1/4$, and $\tau_{pol_{symm}}=\operatorname{asin}(\sqrt{(}1/4))$. Differences are seen between this simplified point-source solution and the solid-black-line complete-form solution, which takes into account the extended disk of the Sun and truncates lines of sight at the observer, as discussed in the text.
@@
Figure 3: Forward modeled $\tau_{pol}$ for $n_{e}=r^{-c}$ spherically symmetric density distributions, for choices of (top) $c=1.9,2.0,2.1$ and (bottom) $c=2.5,3,4$, for elongation angles $\varepsilon<45^{\circ}$.
To establish how generally this slope diagnostic might be used, we need to consider the effects of asymmetric integration limits for lines of sight $\varepsilon>0$. In these cases, we see from Equation 2.3.2 that an additional correction term dependent LOS elongation angle $\varepsilon$ needs to be added:
@@
It is clear that as $\varepsilon\rightarrow 0$, $F_{corr}(\varepsilon)\rightarrow 0$, and similarly as $\varepsilon\rightarrow\frac{\pi}{2}$, $F_{corr}(\varepsilon)\rightarrow 0$. For both of these limits, the symmetric solutions shown in Equations 38-40 are regained.
@@
Something close to a $r^{-2}$ falloff is what we would expect to see for a mass-conserving constant-speed solar wind, so that $\tau_{pol_{r^{-2}}}=\operatorname{asin}(\sqrt{(}1/4))=0.524$ – as can be seen in Figures 2 and 3. Indeed, Figure 4 shows this is the central value in the color bar ranging from minimum to maximum density for a global simulation of a forward-modeled complex solar wind. The associated polarization degree for a $r^{-2}$ falloff is $p=0.6$ (see Equation 38), similar to the asymptotic limit of $0.64$ found in the classical model of Baumbach (1938) (see also Lamy et al. (2020)), and consistent with observations from the ASPIICS spacecraft (Zhukov et al., 2025a).
In summary: White light polarization observations will provide a local diagnostic of density radial falloff for the background solar wind, independent from and complementary to non-local techniques. In addition, we have shown that for observations along lines of sights with elongation angles in the range $\sim 1.5-15^{\circ}$ (the lower part of the PUNCH field of view), the effects of asymmetry are predictable and small enough to distinguish between falloff slopes.